OpenStudy (anonymous):

How to solve for x in this equation 4/x+3/x=1 show steps?

OpenStudy (saifoo.khan):

\[\frac{4+3}{x} = 1\] \[\frac 7x = 1\] \[x = \frac 71\] \[x = 7\]

OpenStudy (anonymous):

Thank you so much! Would you happen to know how this one is done 1/x+1 + 5/x+1=2 is it done the same as that?